# Understanding and Mitigating the Impacts of Differentially Private Census Data on State Level Redistricting

### v1

@@ 6.1 Gingles 1 in the face of MMD discrepancy
Figure 3 illustrates the MMD discrepancies that arise in an ensemble of GA state house plans, which serves well as an illustrative geography due to its significant Black population and large legislature. The histogram breaks down the plans by number of majority-Black MMDs according to the DEMO dataset. Each histogram bin further broken down by MMD discrepancy (i.e. the difference between the number of MMDs as measured by DEMO and SWAP). Across the whole ensemble, the mean MMD discrepancy is 0.77, and 56% of plans have a non-zero MMD discrepancy. Observe that plans with more MMDs have greater discepancies (Pearson’s $r=0.21$). In particular, among plans with the maximum number of MMDs (51), 62.45% of plans have a non-zero discrepancy, which is to say most of the optimized plans contained more MMDs according to DEMO than according to SWAP. As shown in Section 6.2, similar patterns hold in many other state legislative geographies.
@@
Figure 3: MMD discrepancies in the Georgia state house. We plot the number of majority-Black districts in plans sampled in our short bursts ensemble, which is designed to produce plans with many MMDs. Plans are grouped according to the number of majority-Black districts as measured using DEMO. Bars are colored to indicate the fraction of those plans which have the corresponding MMD discrepancy. Each ensemble contains 100,000 plans sampled with a 5% population deviation limit using the DEMO data.
This do not bode well for a plaintiff bringing a vote dilution challenge (under the legal view that the CEF is the appropriate reference). Gingles 1 incentivizes plaintiffs to create demonstration plans with many MMDs. If the resulting plans behave like those in our short bursts ensemble—which optimizes the same quantity—then those plans may be expected to have fewer MMDs under the CEF, depending on the geography.
Rescuing Gingles 1
We argue that the above interpretation misunderstands Gingles 1: demonstration plans can continue to be used as evidence in support of Gingles 1 claims, despite the MMD discrepancies. The relevant legal question is not whether the specific plans adduced have so many MMDs. It is whether they support a conclusion that there exist plans with more MMDs than the enacted plan. (These questions are typically conflated, as they are equivalent when ignoring disclosure avoidance.)
To address this reframed Gingles 1 question, we ask whether the maximum observed number of MMDs within an ensemble is affected by the discrepancy between the two datasets. Concretely, among plans with the most MMDs under DEMO, do any of them have zero MMD discrepancy? The answer is yes — in every short-bursts ensemble we sampled. Our ensembles generated using DEMO always yielded plans that are also maximal when measured using SWAP.
The other side of the Gingles 1 equation is the number of MMDs in the enacted plan. It is conceivable that the enacted plan $\mathcal{P}^{*}$ has more MMDs according to the CEF than according to the Redistricting Data. Let $m$ be the maximum observed number of MMDs in the demonstration plans. A defendant might argue, then, that the enacted plan might maximize the number of MMDs, even if the number measured in the noisy data is less than $m$. This issue would arise only if $\mathsf{MMD}_{\mathsf{demo}}(\mathcal{P}^{*})<m$ and $\mathsf{MMD}_{\mathsf{swap}}(\mathcal{P}^{*})\geq m$. In our ensembles, this is unlikely to happen, as the MMD discrepancy is biased in the other direction. In particular, in 52 of 54 ensembles, the fraction of those plans where $\mathsf{MMD}_{\mathsf{demo}}(\mathcal{P})=m-1$ that also satisfy $\mathsf{MMD}_{\mathsf{swap}}(\mathcal{P})>\mathsf{MMD}_{\mathsf{demo}}(\mathcal{P})$ is less than 1%. In the Massachusetts State House, it is 2.0%, and in the California State House, it is 3.9%.
Taken together, if the Gingles 1 precondition is satisfied in the DEMO data—shown using an illustrative plan—then the precondition is likely satisfied in the SWAP data as well, for our ensembles. This holds even when the illustrative plans are more likely than not to have fewer MMDs in SWAP than in DEMO.
So far, we imagined a plaintiff who has found an illustrative map with more MMDs than an enacted map. Does the 2020 DAS make finding such illustrative maps harder? We consider this question only briefly. For each state legislative geography, we generate one ensemble using DEMO and one using SWAP. For each pair, we compare the observed maxima. They agreed in 42 of the the 54 legislative geographies where any majority-Black districts were observed. Of the remaining 12, the maximum was greater (by one) in the SWAP ensemble in 7 cases. The only geography with a difference greater than one district was the Louisiana House of Representatives, with 37 majority-Black districts in DEMO ensemble and 35 in the SWAP ensemble.
6.2 Exploratory analysis of MMD discrepancies
6.2.1 Discrepancies vary greatly by geography and ensemble
@@
We examine the MMD discrepancies in the 54 geographies where our ensembles found at least one MMD (28 lower house, 26 upper house). Different geographies experience very different levels of MMD discrepancy. Some geographies experience very minor discrepancies; others very significant (e.g., 2.68% of OH upper house plans have non-zero discrepancies, compared to 63.45% of MS lower house plans).
@@
Still, some patterns we see in the GA house generalize to almost all geographies. First, positive MMD discrepancies—more MMDs according to DEMO than SWAP—are both more frequent and larger in magnitude than negative discrepancies. Second, plans with more MMDs have greater discrepancies. Moreover, we observe that MMD discrepancies are generally greater in lower houses than in the (typically smaller) upper houses within the same state.
We briefly compare the MMD discrepancies observed in in our short bursts ensembles to those observed in our base ensembles. The short bursts ensemble generally has greater MMD discrepancies than base ensemble. In the base ensemble, the non-zero discrepancy rate is 9% and the mean discrepancy is 0.057. In the short bursts ensemble, those numbers are 56% and and 0.77, respectively. Even so, the short bursts ensemble always succeeds in finding a plan with more MMDs than the base ensemble in the same geography, even accounting for the greater MMD discrepancies.
6.2.2 Discrepancies appear related to BVAP margin
To try to better understand how MMD discrepancies arise, we briefly look at discrepancies in individual districts in the GA house. We observe that districts with small positive BVAP margins are more likely to experience discrepancies. Note this is a preliminary finding based on just a single geography.
Figure 4: MMD discrepancy rate by BVAP margin for the Georgia state house. We group the distinct districts short bursts ensemble according to the Black voting-age population margin. The height of each bar depicts the district-level MMD discrepancy rate for the group. Namely, the fraction of plans measured as majority Black according to one of DEMO and SWAP, but not both. Districts with small Black majorities in DEMO are much more likely to experience an MMD discrepancy than those with small non-Black majorities.
For each of the individual districts $D$ generated in the GA state house ensembles, we record the BVAP margin: $\mathsf{BVAP}_{\mathsf{demo}}(D)-0.5\cdot\mathsf{VAP}_{\mathsf{demo}}(D)$. Figure 4 groups these districts by BVAP margin, and plots the observed fraction of each group with an MMD discrepancy: majority Black according to one of the DEMO or SWAP, but not both (i.e., $\mathsf{MMD}_{\mathsf{demo}}(D)\neq\mathsf{MMD}_{\mathsf{swap}}(D)$). Note that no district with BVAP margin greater than ±$\pm 150$ had an MMD discrepancy. Unsurprisingly, discrepancies are more common as the BVAP margin approaches zero from either side. More interestingly, discrepancies are much more likely when the BVAP margin is positive—just over 50% Black—than negative. This is consistent with the positive bias in net discrepancies of the plans.
The correlation between small positive BVAP margins and discrepancies helps explain the greater discrepancies observed in the short bursts ensemble than in the base ensemble. (The observed likelihoods of discrepancy by BVAP margin appears similar across the two ensembles.) Districts with small Black majorities are much more common in the short bursts ensemble than the base ensemble (Figure 5 in Appendix B). This is intuitive. Just as packing and cracking produces maps with few MMDs, we would expect maps with many MMDs to look neither packed (large positive margins) or cracked (small negative margins).
@@ 7 Discussion
At the heart of this paper is a legal question. What data should redistricting law treat as the ground truth: the data as enumerated (the CEF) or as published (the Redistricting Data)? We don’t answer the question. Instead, our results suggest that the answer may not greatly affect a redistricter’s ability to comply with the law. While noticeable discrepancies do arise as a result of the new DAS, it is straightforward to account for those discrepancies when reasoning about plans. Though we are not experts in this area of law, we think it is unlikely that disclosure avoidance would cause courts to reject the long acknowledged “legal fiction” that the Decennial Census exactly reflects the underlying population. But others disagree. The case for statistical adjustment has only strengthened with the Census Bureau’s recent release of the so-called Noisy Measurement Files: an intermediate byproduct of the 2020 DAS that can be used to derive statistically better estimates of population than the official Redistricting Data [34].
Our hypothetical data user in this paper is concerned with redistricting plans valid under the CEF. One can instead view the data user as wanting to produce results that would have been valid had the 2010 DAS been used in 2020 (as advocated by critics of the new DAS). For our OPOV analyses, there is no difference between these two views, because the 2010 DAS holds total population invariant. But as the 2010 DAS perturbs race, our analyses of minority representation apply more directly to the latter view, though we believe our overall conclusions to apply in both settings (see Section 3).
7.1 Limitations
An important limitation is that we study algorithmically sampled plans, incorporating only a few redistricting criteria (contiguity, population balance, and compactness). But real districts are drawn by people as part of a political process, taking account of many factors.
Our sample frame does not contain all permissible plans: it excludes plans that subdivide block groups, for instance. Nor does it necessarily contain only permissible plans: ignoring political subdivisions and communities of interest in states where required, for example. It is hard to predict the net effect of these differences our findings. As we explain next, breaking up block groups would tend to strengthen the effects of the noise. On the other hand, the common practice of keeping counties intact would tend to weaken those effects. Future work could use more sophisticated ensemble analyses to account for individual states’ redistricting requirements [21, 35], or perhaps a dataset of hand-drawn plans.
We were limited in our computational resources. In order to widen the geographic scope of our analysis, we chose to use block groups, rather than blocks, as the geounit from which our districts were composed. This greatly reduced the computational complexity, allowing us to run more chains and study larger geographies. But in real-world redistricting, block groups are often split between neighboring districts. To better understand the extent to which this choice impacted our results, we computed the critical offset and no offset discrepancy rate for three small-enough state geographies using blocks as our base geounit (Table 5 in Appendix A). We observe a notable increase in both metrics in all three geographies, indicating that our results may underestimate the impact of the new DAS on real redistricting. We hypothesize the difference is greatest where districts are very small, as in the three examined geographies. This aligns with guidance from the Census Bureau suggesting that data users should aggregate block-level data to increase the accuracy of their analyses.
Another limitation is that comparing the DEMO and SWAP datasets does not isolate the impacts of the 2020 DAS [36]. It reflects the impact of both the 2010 and 2020 DASes. This is inherent. Outside the Census Bureau, it is impossible to directly compare the public 2020 Redistricting Data and the confidential 2020 CEF. DEMO was created to aid stakeholders in understanding the effects of the 2020 DAS by comparing it to SWAP. If we were somehow able to run our analyses using DEMO and the 2010 CEF, nothing in Section 5 would change except possibly in the last paragraph. By using only total population—unaffected by the swapping—we observe the impact of the 2020 DAS alone [37]. On the other hand, Section 6 examines the BVAP as a fraction of the VAP in districts. In 2010, swapping did not affect VAP. But it could affect the BVAP if households with different BVAP were swapped between districts. We cannot correct for this, as the 2010 swapping rates are secret. This is an inherent limitation any work based on comparing the 2010 demo and redistricting data [38].
Other limitations stem from our scope, rather than our methods. Most obviously, we focused on state-legislative redistricting, excluding both Congressional and sub-state redistricting, and on Gingles 1, excluding the other parts of the Gingles framework and racial-gerrymandering litigation more broadly. So while our results give us hope that discrepancies from disclosure avoidance are tolerable, more serious discrepancies may still lurk elsewhere. As a point of comparison, we provide a limited set of results for experiments on Congressional district geographies in Table 6 in Appendix A. Future work could further study Congressional and sub-state redistricting, or extend [9] by quantifying discrepancies the second and third Gingles tests.
Finally, we do not contend with the potential use of the Noisy Measurement File (NMF) for redistricting. While recent work has leveraged the NMF to analyze the error introduced by the 2020 DAS [39], it introduces novel difficulties when applied to redistricting. For example, since hierarchical consistency is not maintained, the population of a district as measured by the NMF can change depending on how its constituent geounits are aggregated. We see addressing this issue and evaluating whether the NMF is legally suitable for this use case as an avenue for future work.
7.2 Additional Related Work
Cohen et al. [9] also use ensemble methods to study the effects of the 2020 DAS on redistricting. Instead of using the Demonstration Data Set, they analyze samples of noised data generated using an early version of the 2020 DAS, albeit with imperfect input data. They conclude that the “the practical effects …[do] not materially threaten any intended uses we considered.” For population balance and Gingles 1, this is based largely on showing the magnitude noise is small relative to the populations considered and to known sources of error. For Gingles 2 and 3, they propose a tweak to the standard method making it more robust to noise. This in part inspired our study of offsets in Section 5.
Among non-ensemble analyses of the impact of the 2020 DAS on redistricting, the most relevant is a study examining the reliability and consistency of measurements of demographic characteristics in geographies of different sizes [40]. These are important for VRA enforcement, but don’t do away with issues presented by the sharp legal thresholds we consider. Other works study the impact of the 2020 DAS on different policy issues, including minority representation [41], misallocation of federal funds [42], and public health monitoring [43], for example.

### v2

@@ 6.2.1 Georgia State House Plans
Figure 5 illustrates the MMD discrepancies that arise in 100,000 plans sampled using our base merge-split MCMC algorithm (Section 3.2). The histogram breaks down the plans by number of majority-Black MMDs according to the DEMO dataset. Each histogram bin is further broken down by MMD discrepancy: the difference between the number of MMDs as measured by DEMO and SWAP. Across the whole ensemble, the mean MMD discrepancy is 0.03, and 9% of plans have a non-zero MMD discrepancy. Furthermore, plans with more MMDs have slightly greater discrepancies (Pearson’s $r=0.08$).
@@
Figure 5: MMD discrepancies in the Georgia state house base ensemble. We plot the number of majority-Black districts in plans sampled in our base ensemble. Plans are grouped according to the number of majority-Black districts as measured using DEMO. Bars are colored to indicate the fraction of those plans which have the corresponding MMD discrepancy. Each ensemble contains 100,000 plans sampled with a 4.4% population deviation limit (utilizing $\Delta=0.6\%$) using the DEMO data.
Figure 6 illustrates the results of the same analysis for 5,000 plans generated using short-bursts optimization These numbers are markedly higher: the mean MMD discrepancy is 1.02, and 66% of plans have a non-zero MMD discrepancy. The correlation between number of MMDs and MMD discrepancy is also higher (Pearson’s $r=0.16$).
Figure 6: MMD discrepancies in the Georgia state house optimized ensemble. We plot the number of majority-Black districts in plans sampled in our short bursts ensemble, which is designed to produce plans with many MMDs. Plans are grouped according to the number of majority-Black districts as measured using DEMO. Bars are colored to indicate the fraction of those plans which have the corresponding MMD discrepancy. Each ensemble contains 100,000 plans sampled with a 4.4% population deviation limit (utilizing $\Delta=0.6\%$ offset listed in Table 4) using the DEMO data.
6.2.2 Other State Legislative Geographies
@@
We examine the MMD discrepancies in the 52 geographies where our short burst ensembles found at least one MMD (27 lower house, 25 upper house). Full results are included in Table 8 in Appendix A. Different geographies experience very different levels of MMD discrepancy. Some geographies experience very minor discrepancies; others very significant (e.g., 0.9% of CA lower house plans have non-zero discrepancies, compared to 78.9% of MS lower house plans).
@@
Still, some patterns we see in the GA house generalize to almost all geographies. First, positive MMD discrepancies—more MMDs according to DEMO than SWAP—are both more frequent and larger in magnitude than negative discrepancies. Second, plans with more MMDs have greater discrepancies. Finally, we observe that MMD discrepancies are generally greater in lower houses than in the (typically smaller) upper houses within the same state.
6.2.3 DEMO vs SWAP vs CEF
[…]
Still, characterizing the expected discrepancies between DEMO and SWAP can be useful for redistricters in practice. If one accepts the need for disclosure avoidance in census tabulations but is concerned that the new DAS yields data that is less fit for use than the previous DAS, our results should provide some reassurance. Our results can be seen as useful for understanding whether DEMO is sufficiently close to SWAP for use in redistricting, setting aside whether SWAP is sufficiently close to the 2010 CEF to be fit for use. If it is, and if SWAP was fit for use for redistricting (which was never under question), then that lends support for DEMO’s fitness for use as well.
6.3 The Connection between BVAP Margin and MMD Discrepancy
Above, when drawing districts with DEMO and measuring against SWAP we see: (1) a bias towards more MMDs in DEMO than SWAP, and (2) the bias increases as the number of MMDs increases within a geography. This section describes a simple probabilistic model that helps explain these observations. We do so by connecting plan-level discrepancies to a district-level quantity (BVAP margin) whose sign determines whether a district is counted as majority-Black.
BVAP margin
The key quantity is BVAP margin: the size of a district’s Black voting age majority (i.e., Black voting age population minus half of the total voting age population). We denote a district’s BVAP margins under DEMO and SWAP as:
$B_{\mathsf{demo}}(D)$ $=\mathsf{BVAP}_{\mathsf{demo}}(D)-0.5\cdot\mathsf{VAP}_{\mathsf{demo}}(D)$ (1)
$B_{\mathsf{swap}}(D)$ $=\mathsf{BVAP}_{\mathsf{swap}}(D)-0.5\cdot\mathsf{VAP}_{\mathsf{demo}}(D)$
By definition, a district is an MMD according to DEMO if and only if its BVAP margin under DEMO is positive: $B_{\mathsf{demo}}(D)>0$. Likewise, for SWAP.
Given only the DEMO dataset, a redistricter can reason about MMDs under SWAP by observing the following:
$B_{\mathsf{swap}}(D)=B_{\mathsf{demo}}(D)+\underbrace{\bigl(B_{\mathsf{swap}}(D)-B_{\mathsf{demo}}(D)\bigr)}_{\mathsf{err}_{\mathsf{das}}^{\mathsf{MMD}}(D)}$ (2)
This decomposition separates what is observable by the redistricter from what is not (as Section 5.2 did for population balance). The first component $B_{\mathsf{demo}}(D)$ is entirely under the control of the mapmaker. The second component is the change in BVAP margin from disclosure avoidance, which we denote $\mathsf{err}_{\mathsf{das}}^{\mathsf{MMD}}(D)$.
By equation (2), a district contributes to plan-level MMD discrepancy precisely when the disclosure-avoidance term $\mathsf{err}_{\mathsf{das}}^{\mathsf{MMD}}(D)$ causes a sign flip between $B_{\mathsf{demo}}(D)$ and $B_{\mathsf{swap}}(D)$. Districts with $B_{\mathsf{demo}}(D)\approx 0$ are the most likely to flip. Therefore, we might hope to explain characteristics of the plan-level MMD discrepancies by looking at the characteristics of $B_{\mathsf{demo}}(D)$ and $\mathsf{err}_{\mathsf{das}}^{\mathsf{MMD}}(D)$, and focusing specifically on near-threshold districts where $B_{\mathsf{demo}}(D)\approx 0$.
Understanding biases in MMD discrepancy
In our experimental setup, we tend to see a bias towards more MMDs in DEMO than in SWAP. This corresponds to plans having more districts $D$ satisfying $B_{\mathsf{swap}}(D)<0<B_{\mathsf{demo}}(D)$ than the reverse. The more MMDs a plan has, the greater the observed bias tends to be. Why might this be?
The reasoning above suggests that we find an explanation by looking at biases in $B_{\mathsf{demo}}(D)$ and $\mathsf{err}_{\mathsf{das}}^{\mathsf{MMD}}(D)$, focusing on near-threshold districts where the BVAP margin is close to 0 (i.e., where sign flips / discrepancies are plausible).
Hence, we consider two possible sources for the observed bias. First, if $\mathsf{err}_{\mathsf{das}}^{\mathsf{MMD}}(D)$ is negatively biased. Second, if plans tend to have more districts $D$ with BVAP just over 50% ($B_{\mathsf{demo}}(D)>0$) than just under 50% ($B_{\mathsf{demo}}(D)<0$). The latter would also help explain the correlation between the number of MMDs and the bias, as maximizing the number of MMDs requires smaller Black majorities on average.
Both explanations are borne out in our data for the Georgia state house. Figure 7 plots $\mathsf{err}_{\mathsf{das}}^{\mathsf{MMD}}(D)$ for the districts in the short-burst ensemble. For districts with BVAP margin close to 0, the mean $\mathsf{err}_{\mathsf{das}}^{\mathsf{MMD}}(D)$ is 22.303-22.303 (standard deviation 6.3). [^13] Across all districts, the mean $\mathsf{err}_{\mathsf{das}}^{\mathsf{MMD}}(D)$ is $0.030$.
Figure 7: Empirical change in BVAP margin. Change in BVAP margin ($\mathsf{err}_{\mathsf{das}}^{\mathsf{MMD}}(D)=B_{\mathsf{swap}}(D)-B_{\mathsf{demo}}(D)$) among districts in our short bursts ensemble for the Georgia State House of Representatives. The blue histogram includes all districts that were contained within plans in the ensemble. The orange histogram is restricted to districts with Black voting age populations within 196 individuals of half of the total VAP. The threshold of 196 was the maximum error in BVAP margin observed in the ensemble. Note that the noise in BVAP margin for districts that have close to 50% BVAP is biased in the negative direction.
Figure 9 plots histograms of $B_{\mathsf{demo}}(D)$ for all districts in both the base and short-burst ensembles. In the short bursts ensemble, districts with small Black majorities are much more common than districts with small non-Black majorities. This is not true in the base ensemble which ignores race, helping explain the smaller biases seen in those plans.
A model of MMD discrepancies from BVAP margins
The discussion above suggests a simple probabilistic model for MMD discrepancies using only BVAP margins, ignoring all other effects (e.g., correlations between errors across districts). The model aims to capture the behavior of near-threshold districts whose BVAP is close enough to 50% that disclosure-avoidance error could change their majority status.
Given a plan with districts $D_{i}$, sample $B_{\mathsf{swap}}(D_{i})=B_{\mathsf{demo}}(D_{i})+E_{i}$, where $E_{i}\sim\mathrm{N}(\mu,\sigma^{2})$ is normally distributed. The parameters $\mu$ and $\sigma$ are estimated from the ensemble by looking at districts whose BVAP margins are small enough that a sign flip is possible (i.e., $|B_{\mathsf{demo}}(D)|\leq\max_{D^{\prime}}|\mathsf{err}_{\mathsf{das}}^{\mathsf{MMD}}(D^{\prime})|$, the largest observed change in BVAP margin). Intuitively, this conditions on the at-risk districts for which noise can change MMD classification, and avoids letting the many districts with very large margins dominate the fit.
Figure 8 plots the results of applying this simulated MMD discrepancy model to the 5,000 Georgia state house plans in our short-bursts ensemble ($\mu=-22.303$, $\sigma^{2}=39.514$), side-by-side with the MMD discrepancies observed when comparing DEMO and SWAP. The two plots contain exactly the same plans each with the same number of MMDs in DEMO. The only difference is whether MMDs in SWAP are measured against the real data or sampled from the model above.
Figure 8: Modeling MMD discrepancies in the Georgia state house. We compare the MMD discrepancies previously observed in Figure 6 to the MMD discrepancies predicted by our synthetic noise model for the Georgia state house.
Qualitatively, the model appears to explain much of the observed biases in MMD discrepancies, but not all. The model is deliberately simple and does not capture all features of the observed discrepancies (e.g., non-normality and correlations in $\mathsf{err}_{\mathsf{das}}^{\mathsf{MMD}}(D)$), so it should be interpreted as a clarifying approximation rather than a full predictive account. More work would be needed to quantify the result and to understand what characteristics of the data are driving the remaining bias.
Offsets for Gingles 1 plans
As in Section 5, a natural approach to mitigating MMD discrepancies in the context of Gingles 1 would be to draw illustrative plans whose majority-Black districts have large BVAP margins. Instead of targeting 50% VAP + 1 person, target 50% + 1 + offset, for some value offset>0>0.
We leave a detailed study of the effectiveness of offsets for MMD discrepancies to future work. For now, we remark that this is already common practice. For example, across the eleven illustrative plans used in the recent Supreme Court case Allen v. Milligan (2023), the smallest BVAP margin in a majority-Black district was 256 people (50.05% BVAP) [34], larger than any $\mathsf{err}_{\mathsf{das}}^{\mathsf{MMD}}(D)$ observed in the Georgia ensemble.
7 Limitations
Methods
An important limitation is that we study algorithmically sampled plans, incorporating only a few redistricting criteria (contiguity, population balance, and compactness). But real districts are drawn by people as part of a political process, taking account of many factors. One should not read too much into the specific numbers we observe: changing the way we generate districts would likely change the quantitative results.
Some algorithmic methods for sampling plans have the ability to generate representative samples of districting plans under relevant probability distributions for redistricting [31, 2]. For the purposes of this paper, we make no such claims and ignore many of the complexities of practical redistricting. At the same time, from a technical point of view, we believe it is reasonable to offer statistics over a set of interest as a descriptive matter without a need for a distributional claim about the sample. One goal of this paper is to replicate and extend findings from prior work, and to explain the phenomena we observe. For such uses, having a clear, well-defined sample frame may be more useful than strict adherence to human redistricting practice.
Our sample frame does not contain all permissible plans: it excludes plans that subdivide precincts, for instance. Nor does it necessarily contain only permissible plans: ignoring political subdivisions and communities of interest in states where required, for example.
It is hard to predict the net effect of these differences on our findings. As we explain next, breaking up precincts would tend to strengthen the effects of the noise. On the other hand, the common practice of keeping cities and counties intact would tend to weaken those effects. Future work could use more sophisticated ensemble analyses to account for individual states’ redistricting requirements [6, 32], or perhaps a dataset of hand-drawn plans.
In order to make the most of our limited computational resources, we chose to use precincts as the geounit from which our districts were composed, rather than blocks. This greatly reduced the computational complexity, allowing us to run more chains and study larger geographies. In practice, redistricters may split precincts between neighboring districts (though they often prefer not to).
To get some indication as to how using blocks would have affected our results, we computed the critical offset and no offset discrepancy rate for three small-enough state geographies using blocks as our base geounit (Table 5 in Appendix A). We observe a notable increase in both metrics in all three geographies, indicating that our results may underestimate the impact of the new DAS on redistricting in practice. (On the other hand, keeping cities and counties whole would have the opposite effect.) We hypothesize the difference is greatest where districts are very small, as in the three examined geographies. This aligns with guidance from the Census Bureau suggesting that data users should aggregate block-level data to increase the accuracy of their analyses.
Data
Comparing the DEMO and SWAP datasets does not isolate the impacts of the 2020 DAS [7]. It reflects the impact of both the 2010 and 2020 DASes. This is inherent. Outside the Census Bureau, it is impossible to directly compare the public 2020 Redistricting Data and the confidential 2020 CEF. DEMO was created to aid stakeholders in understanding the effects of the 2020 DAS by comparing it to SWAP.
If we were somehow able to run our analyses using DEMO and the 2010 CEF, nothing in Section 5 would change except possibly in the last paragraph. By using only total population—unaffected by the swapping—we observe the impact of the 2020 DAS alone [26].
On the other hand, Section 6 examines the BVAP as a fraction of the VAP in districts. In 2010, swapping did not affect VAP. But it could affect the BVAP if households with different BVAP were swapped between districts. This is an inherent limitation any work based on comparing the 2010 demonstration and redistricting data [40].
One way to try to isolate the effect of the 2020 DAS would be to run it ourselves using an alternate dataset (e.g., reconstructed 2010 Census microdata derived from SWAP as in [12]). But running the 2020 DAS reliably is difficult [4] and beyond the scope of the present work.
In Appendix B, we instead take a step towards isolating the effect of the 2010 DAS, using data provided to us by the authors of [4]. More work would be needed to fully isolate the effects of the two disclosure avoidance systems. Still, based on our analysis in Section 6.3 and Appendix B, we believe that any differences in the MMD discrepancies arising from the two systems are more likely differences in degree than in kind.
Scope
Other limitations stem from our scope, rather than our methods. Most obviously, we focused on state legislative redistricting, excluding both Congressional and sub-state redistricting, and on Gingles 1, excluding the other parts of the Gingles framework and racial-gerrymandering litigation more broadly. So while our results give us hope that discrepancies from disclosure avoidance are tolerable, more serious discrepancies may still lurk elsewhere. As a point of comparison, we provide a limited set of results for experiments on Congressional district geographies in Table 6 in Appendix A. Future work could further study Congressional and sub-state redistricting, or extend [12] by quantifying discrepancies for the second and third Gingles tests.
Finally, we do not contend with the potential use of the Noisy Measurement Files (NMF) for redistricting (see below). While recent work has leveraged the NMF to analyze the error introduced by the 2020 DAS [25], it introduces novel difficulties when applied to redistricting. For example, since hierarchical consistency is not maintained, the population of a district as measured by the NMF can change depending on how its constituent geounits are aggregated. Addressing this issue and evaluating whether the NMF is suitable for redistricting uses is an avenue for future work.
@@ 8 Discussion
In its lawsuit challenging the 2020 DAS, Alabama argued that One-Person, One-Vote requires balancing populations according to the confidential CEF data, rather than the official Redistricting Data (Alabama v. Department of Commerce, 2021). In the face of discrepancies, Alabama concluded that redistricting plans created using the Redistricting Data could be illegal. As Alabama put it, this could potentially “inhibit” or “impede” the redistricting process (Alabama v. Department of Commerce, 2021).
What data should redistricting law treat as the ground truth: the data as collected, edited, and imputed (the CEF) or as published (the Redistricting Data)? Our experiments can’t answer the question. Instead, they suggest that treating the CEF as the ground truth may not greatly affect a redistricter’s ability to comply with the law using only the Redistricting data. At the very least, there exist countermeasures for our specific redistricting procedure. To the extent that the finding generalizes, data discrepancies may “impede” redistricters by giving them something else to account for, but appears unlikely to “inhibit” lawful redistricting altogether.
At least for state legislative redistricting, it appears that one can use the Redistricting Data to meet OPOV and MMD threshold requirements on the unseen data. Discrepancies do arise, but can be understood and accounted for. For OPOV requirements, using offsets does not make generating state legislative districts more difficult. In particular, all our ensembles in Section 6 used the critical offsets found in Section 5. And for counting MMDs, the common practice for producing illustrative maps is already to avoid very small margins.
We caution against over-interpreting our results as saying that disclosure avoidance can have no practical impacts on redistricting. Prior work highlights many plausible interactions [24, 12, 48], only a few of which we explore. The perturbations from disclosure avoidance (or other sources of error) could conceivably make it harder to challenge a racially gerrymandered plan if, for example, a cracked plan in a racially polarized state appears to have more majority-minority districts in the DEMO data than it has “effective majority” districts in reality. Perhaps our probabilistic model of MMD discrepancies might help us reason about this possibility, but we leave that to future work.
Though we are not experts in this area of law, we think it is unlikely that changes in disclosure avoidance would cause courts to reject the long acknowledged “legal fiction” that the Decennial Census exactly reflects the underlying population. On the other hand, the case for statistical adjustment has only strengthened with the Census Bureau’s recent release of the so-called Noisy Measurement File (NMF) [41]. These are the raw, noisy statistics produced by the 2020 DAS before being post-processed into the required tabular form. By definition, the NMF is statistically more informative than any derived data products, including the Redistricting Data itself. It is conceivable that NMF-based statistical adjustments will find their place redistricting or policymaking. For the moment, this is mere speculation. The official Redistricting Data remains the only basis for redistricting today, to the best of our knowledge.